Q11: What is the value of a stock expected to pay a constant $5 dividend each year forever, if the market required
rate of return is 18%?

Q12: A stock just paid an annual dividend of $2. The dividends are expected to grow at 20% per year over each of
the next three years and 5% per year thereafter. What is the value of the stock if the required rate of return is
12%?
Answer #1

1.Current price=D1/(Required return-Growth rate)

=5/0.18

=$27.78(Approx).

2.

D1=(2*1.2)=2.4

D2=(2.4*1.2)=2.88

D3=(2.88*1.2)=3.456

Value after year 3=(D3*Growth rate)/(Required return-Growth rate)

=(3.456*1.05)/(0.12-0.05)=$51.84

Hence value of the stock=Future dividends*Present value of discounting factor(rate%,time period)

=2.4/1.12+2.88/1.12^2+3.456/1.12^3+51.84/1.12^3

which is equal to

=$43.80(Approx).